OpenStudy (anonymous):

sqrt{x^2-4}=2sqrt{3}

OpenStudy (anonymous):

\[\sqrt{x^2-4}=2\sqrt{3}\]

OpenStudy (anonymous):

x^2-4=12 x^2=16 x=+or-4